# Continuous-time persuasion by filtering

### v2

@@ Abstract
We frame dynamic persuasion in a partial observation stochastic control Leader-Follower game with an ergodic criterion. The Receiver controls the dynamics of a multidimensional unobserved state process. Information is provided to the Receiver through a device designed by the Sender that generates the observation process. The commitment of the Sender is enforced. We develop this approach in the case where all dynamics are linear and the preferences of the Receiver are linear-quadratic. We prove a verification theorem for the existence and uniqueness of the solution of the HJB equation satisfied by the Receiver’s value function. An extension to the case of persuasion of a mean field of interacting Receivers is also provided. We illustrate this approach in two applications: the provision of information to electricity consumers with a smart meter designed by an electricity producer; the information provided by carbon footprint accounting rules to companies engaged in a best-in-class emissions reduction effort. In the first application, we link the benefits of information provision to the mispricing of electricity production. In the latter, we show that information provision is a substitute for best-in-class objective.
@@ 1 Introduction
We find that emulation between companies tends to lower the need for information provision in the form of constraining accounting rules. In fact, our analysis shows that information provision and best-in-class objective target by firms are substitute. Both factors tend to reduce the average carbon footprint, and thus, a sector engaged in an emulation process for carbon footprint reduction reduces the need for stringent accounting rules.
@@ 3 Linear-quadratic ergodic persuasion game
Let us consider a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$ endowed with two independent Brownian motions, $W$ and $B$, of dimensions $d_{W}$ and $d_{B}$. We focus on a game between a Sender and a Receiver of information, also called the agents, who are indicated, respectively, by S and R. Borrowing the terminology from stochastic filtering theory, the $\mathbb{R}^{d_{W}}$-valued stochastic process $X=(X_{t})_{t\geq 0}$ is called unobservable signal, since the Receiver has no access to it, while he has access to the $\mathbb{R}^{d_{B}}$-valued process $M=(M_{t})_{t\geq 0}$. The latter represents the message that the Sender is providing to him. We assume the following linear dynamics for the state variables:
@@
We say that any $\mathbb{R}^{r}$-valued stochastic process $v$ is admissible if it belongs to $\mathcal{H}_{r}^{2}(\mathbb{F}^{\rm R})$ and it satisfies
@@
The set of all such processes is denoted by $\mathcal{A}^{\rm R}$.
@@ 4.2 Carbon footprint accounting rules
$\sigma^{\ast}_{\infty}:=\frac{1}{2}\Bigg{[}\sqrt{\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}^{2}+4\frac{\delta-\epsilon m_{0}m_{1}\underline{\sigma}}{1+\epsilon^{2}m_{1}^{2}}}+\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Bigg{]},\quad\text{with}\quad\delta:=\frac{\eta}{c\Delta\sigma}.$
(a) (b) (c)
Refer to caption Refer to caption Refer to caption
(d) (e) (f)
Refer to caption Refer to caption Refer to caption
@@
Figure 2: As a function of $\epsilon$ (a) $m^{\ast}_{\infty}$ (b) $\sigma^{\ast}_{\infty}$ (c) Damage cost with optimal information provision and without any information provision. Parameters: $\ell=a=1$, $\gamma=0.1$, $c=0.5$, $\lambda_{a}=0.75$, $\lambda_{q}=0.25$; $\kappa=0.05$ for (a), (b) and (c); $\kappa=0.09$ for (d), (e) and (f).
Figure 2 (a), (b), (c) and (d) display $m_{\infty}^{\ast}$ and $\sigma_{\infty}^{\ast}$ as a function of the best-in-class parameter $\epsilon$ in a situation where the firm puts more weight on its own personal target rather than the target of going close to the best-in-class target $q^{\ast}$. At equilibrium, only the product $\epsilon\lambda_{q}$ matters to measure the intensity with which the representative firm takes into account the best-in-class target. The main observation to be made on the information provision process is that it is a substitute to best-in-class objective. Indeed, we observe first that in (b), $\sigma_{\infty}^{\ast}$ increases with $\epsilon$, meaning that the higher the best-in-class objective intensity, the less the regulator is prone to impose stringent accounting rules. This phenomenon appears more clearly when the mean-reversion coefficient $\kappa$ is increased. In (a) and (b) its value is $\kappa=0.05$ while it is increased to $\kappa=0.09$ in (d) and (e). Increasing $\kappa$ leads to increasing the speed at which firms can identify the true value of their carbon footprint. We observe then that for large values of $\epsilon$, the information provision is no longer needed. This observation translates in terms of costs. Figure 2 (c) shows the decomposition of the cost function of the Sender. The black dotted line represents the cost in the absence of information provision, and thus, the damage. Providing information (figure (c) and (f)), always improves the objective function of the regulator whatever the best-in-class intensity $\epsilon$. However, the gain decreases with $\epsilon$ and might disappear in case of large mean-reversion values.
As a result, if we interpret the information provision as the level of stringency of the accounting rules, our model of information provision shows that the rules are a substitute for best-in-class objective target of a set of firms.

### v3

@@ Abstract
We frame dynamic persuasion in a partial observation stochastic control Leader-Follower game with an ergodic criterion. The Receiver controls the dynamics of a multidimensional unobserved state process. Information is provided to the Receiver through a device designed by the Sender that generates the observation process. The commitment of the Sender is enforced. We develop this approach in the case where all dynamics are linear and the preferences of the Receiver are linear-quadratic. We prove a verification theorem for the existence and uniqueness of the solution of the HJB equation satisfied by the Receiver’s value function. An extension to the case of persuasion of a mean field of interacting Receivers is also provided. We illustrate this approach in two applications: the provision of information to electricity consumers with a smart meter designed by an electricity producer; the information provided by carbon footprint accounting rules to companies engaged in a best-in-class emissions reduction effort. In the first application, we link the benefits of information provision to the mispricing of electricity production. In the latter, we show that even in the absence of information cost, it might be optimal for the regulator to blur information available to firms to prevent them from coordinating on a higher level of carbon footprint to reduce their cost of reaching a below average emission target.
@@ 1 Introduction
Our model of information provision shows that when there is no strategic interaction between firms, only information cost can justify deviation from sending full information. But, in the presence of strategic interaction in the form of a best-in-class emulation process, information can also be used by strategic firms to coordinate on a collectively higher level of carbon footprint to look individually cleaner. And, thus, even in the absence of information cost, it may be optimal for the regulator to blur information available to firms.
@@ 3 Linear-quadratic ergodic persuasion game
Let us consider a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$ endowed with two independent Brownian motions, $W$ and $B$, of dimensions $d_{W}$ and $d_{B}$. We denote by $\mathbb{F}^{W,B}=(\mathcal{F}_{t}^{B,W})_{t\geq 0}$ the filtration generated by $B,W$ and completed with the $\mathbb{P}$-negligible events in $\mathcal{F}$, so fulfilling the usual conditions of right-continuity and $\mathbb{P}$-completeness.
We focus on a game between a Sender and a Receiver of information, also called the agents, who are indicated, respectively, by S and R. Borrowing the terminology from stochastic filtering theory, the $\mathbb{R}^{d_{W}}$-valued stochastic process $X=(X_{t})_{t\geq 0}$ is called unobservable signal, since the Receiver has no access to it, while he has access to the $\mathbb{R}^{d_{B}}$-valued process $M=(M_{t})_{t\geq 0}$. The latter represents the message that the Sender is providing to him. We assume the following linear dynamics for the state variables:
@@
We say that any $\mathbb{R}^{r}$-valued stochastic process $v$ is admissible if it belongs to $\mathcal{H}_{r}^{2}(\mathbb{F}^{B,W})$, it is $\mathbb{F}^{\rm R}$-progressively measurable and it satisfies
@@
The set of admissible processes is denoted by $\mathcal{A}^{\rm R}$.
Remark 3.4.
Observe that the filtration $\mathbb{F}^{\rm R}$ depends implicitly on the control process $v$, that we omitted from the notation for simplicity. Indeed, $v$ affects $X$, which in turns affects $M$, which generates $\mathbb{F}^{\rm R}$. So, defining the Receiver’s admissible controls as simply those processes which are $\mathbb{F}^{\rm R}$-progressively measurable would have resulted in some circularity. The definition above addresses this issue as follows: first, given any $v\in\mathcal{H}_{r}^{2}(\mathbb{F}^{B,W})$, the state variable $(X,M)$ as well as the filtration $\mathbb{F}^{\rm R}$ are well defined and they all depend on $v$; second, we have restricted Receiver’s choice to those controls in $\mathcal{H}_{r}^{2}(\mathbb{F}^{B,W})$ which are also $\mathbb{F}^{\rm R}$-progressive measurable and satisfy the limiting condition above. This procedure is pretty standard in stochastic control with partial observation, see, e.g., the recent paper Cohen et al. (2025) and the discussion therein.
@@ 4.2 Carbon footprint accounting rules
$\sigma^{\ast}_{\infty}:=\frac{1}{2}\Bigg{[}\sqrt{\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}^{2}+4\frac{\delta-\epsilon m_{0}m_{1}\underline{\sigma}}{1+\epsilon^{2}m_{1}^{2}}}+\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Bigg{]},\quad\text{with}\quad\delta:=\frac{\eta}{c\Delta\sigma}.$ (4.15)
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
@@
Figure 2: Case without information cost. As a function of the weight $\lambda_{q}$ of the best-in-class target (a) Optimal choice of standard deviation $\sigma^{\ast}_{\infty}$ of $X^{\ast}_{\infty}$ (b) Optimal $m^{\ast}_{\infty}$ of $X^{\ast}_{\infty}$ (c) Optimal value of the cost function of the Regulator (d) Scaled nformation map ie ratio $(\overline{\sigma}-\sigma^{\ast}_{\infty})/\Delta\sigma$ as a function of both $\lambda_{q}$ and $\epsilon$. Parameters: $\ell=a=10$, $\gamma=1$, $c=0.5$, $\kappa=0.12$, $\epsilon=0.5$.
At this stage, some remarks can be done. First, note that , for a given level of information as measured by the standard deviation $\sigma^{\ast}_{\infty}$ of the carbon footprint process, the expected carbon footprint of the representative firm is a decreasing function of the best-in-class intensity parameter $\epsilon$. Thus, the more the firms are engaged in an emulation process in reducing their carbon footprint, the more the average carbon footprint does reduce. This is the phenomenom that information provision is supposed to reinforce or at least to help in its implementation. But, we already observed that $\sigma^{\ast}_{\infty}(b)$ is a decreasing function of the accounting norm stringency as measured by the regulator’s control $b$. Thus, when the regulator provides more information to firms engaged in a best-in-class emulation process in carbon footprint reduction, the first effet of information provision is to increase the average carbon footprint of the representative firm. Indeed, information helps firms to coordinate in a process to reduce the effort of reaching a below average carbon footprint.
Second, note that the objective of the regulator is to miminise $\mathbb{E}[(X_{\infty}^{\ast})^{2}]=(m_{0}-\epsilon m_{1}\sigma^{\ast}_{\infty})^{2}+(\sigma^{\ast}_{\infty})^{2}$ where $\sigma^{\ast}_{\infty}\in(\underline{\sigma},\overline{\sigma}]$. Thus, when $\epsilon>0$, the regulator has to deal with a tradeoff where information reduces the variance $(\sigma_{\infty}^{\ast})^{2}$ of the firms carbon footprint but increases its average $m_{\infty}^{\ast}$. When, there is no best-in-class target (ie $\epsilon=0$ or $\lambda_{q}=0$), only the existence of an information cost function $G(z)$ prevents the regulator from sending full information and taking $\sigma_{\infty}^{\ast}=\underline{\sigma}$. This point is obtained directly in relation (4.15) by taking $\epsilon=0$. But as soon as firms are engaged in such an emulation process, even in the absence of information cost, the regulator may choose to deviate from sending full information on the state process. The Figure 2 illustrates this phenomenon.
Figure 2 (a), (b), (c) and (d) display $\sigma_{\infty}^{\ast}$, $m_{\infty}^{\ast}$ and the cost function of the regulator as a function of the best-in-class parameter $\epsilon$ in a situation where the is no information cost ($\eta=0$). We observe in picture (a) that for low values of $\epsilon$, it is optimal to send full information: $\sigma^{\ast}_{\infty}$ is set at its lowest possible value. But as soon as $\epsilon$ excesses a thresold, the regulator restrains information. As $\epsilon$ increases, the regulator decreases information until it reaches the largest possible value of $\sigma^{\ast}_{\infty}$. At this stage, we observe in pictures (b) and (c), that the average $m^{\ast}_{\infty}$ has reached its minimum possible value and the regulator’s cost is equal to the case of no information provision, which is also the largest possible cost she can get.
But, more spectacular, we observe that when $\epsilon$ continues to increase, it reaches a second threshold where it is now again optimal to send information. Indeed, picture (b) shows that if the regulator would keep sending no information, the average value $m^{\ast}_{\infty}$ would become highly negative. But, because the regulator’s criterion penalises also large negative values (as if she would not believe in carbon compensation), providing information helps the firms to maintain their average carbon fooprint close to zero.
The picture (d) of Figure 2 provides the information map of the regulator, ie the ratio $(\overline{\sigma}-\sigma^{\ast}_{\infty})/\Delta\sigma\in[0,1]$ as a function of both $\epsilon$ and $\lambda_{q}$. The blue zones correspond to a ratio of $1$ while the dark red zones correspond to a ratio of 00. It shows that the set of points where it is optimal for the regulator to send full information on the firms carbon footprint are not connex. They are separated by an information dessert where information is detrimental to a reduction of carbon footprint.
As a result, our model shows that only information cost can justify deviation from sending full information when there is no strategic interaction between firm willing to reduce their carbon footprint. But, in the presence of strategic interaction in the form of a best-in-class emulation process, information can also be used by strategic firms to coordinate on a collectively higher level of carbon footprint to look individually cleaner. And, thus, even in the absence of information cost, it may be optimal to blur information available to firms.